a ball of mass 100g is thrown vertically from the edge of a table with a speed of 1.3 m/s to compress a spring of a coefficient of 120 n/m. what is the height of the table if the spring compressed to 10.6 cm


1
Expert's answer
2021-08-11T09:54:00-0400

Mass of ball, "m=100\space g=100\times10^{-3}\space kg=10^{-1}\space kg"

Speed of ball, "v=1.3\space m/s"

Spring constant, "k=120\space N/m"

Compression in spring, "x=10.6\space cm=0.106\space m"

Let height of the table "= h"

"\dfrac{1}{2}mv^2+mgh=\dfrac{1}{2}kx^2"

"h=\dfrac{(kx^2-mv^2)}{2mg}=0.601\space m"